OpenStudy (sasogeek):

#fanday friday: IFF ½ of 8 = 5 and the same proportion holds, what is the value of 1/5 of 24?

OpenStudy (kinggeorge):

I'm gonna say.... 6.

OpenStudy (kinggeorge):

If \(8/2=5\), then \(8/5=2\). Since 24=8*3, \(24/5=3*2=6\).

OpenStudy (sasogeek):

okay so that was very fast and you're right lol :) but I'm ur fan already so.... :P but cool xD

OpenStudy (ddcamp):

Is this basically it?